How many different signals can be made by using at least 4 different flags if there are 6 different flags from which to select?

Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)
Answer by greenestamps(13200)



Assuming a "signal" is a specific ordered sequence of flags in a single row....

4 flags: 6*5*4*3 = 360
5 flags: 6*5*4*3*2 = 720
6 flags: 6*5*4*3*2*1 = 720

Add....